OpenStudy (anonymous):

Find the dimensions of the rectangle given that the area is 504 sq. units. the side lengths of the rectangle are: 20x-48 and 7

OpenStudy (anonymous):

A=l*w solve for x, then insert the value back into the side length.

OpenStudy (anonymous):

so to set it up, it would like like: (20x-48)(7) = 504 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

or 504units^2

OpenStudy (anonymous):

No square..

OpenStudy (anonymous):

\((20x-48) \times(7) = 504\) Solve this one only.